The perimeter of a rectangle is 168 inches. the length exceeds the width by 6 inches. Find the length and width.
Answer by mananth(16949)

let the width be x
The length will be x+6
Perimeter = 2L + 2W
= 2(x+6)+2x
2(x+6)+2x=168
2x+12 +2x=168
4x=168-12
4x=156
x=39 the width
Length = 45 inches
39+39+45+45=168